OpenStudy (anonymous):

-9x + 4y = 8 -3x − y = 4?

OpenStudy (kamibug):

-3x-y=4 -y=4+3x y=-4-3x -9x+4(-4-3x)=8 -9x-16-12x=8 -21x=24 x=-24/21 x=-8/7 -3(-24/21) - y = 4 72/21 - y = 4 -y = 4 - 72/21 -y=12/21 y = -4/7